OpenStudy (anonymous):

∫ X^3/(X^2+4)^2dx ∫ X^3/(X^2+4)^2dx @Mathematics

OpenStudy (agreene):

\[\int \frac{x^3}{(x^2+4)^2}dx\] yes?

myininaya (myininaya):

\[\frac{x^3}{(x^2+4)^2}=\frac{Ax+B}{x^2+4}+\frac{Cx+D}{(x^2+4)^2}=\frac{(Ax+B)(x^2+4)+(Cx+D)}{(x^2+4)^2}\]

myininaya (myininaya):

=> \[x^3=Ax^3+4Ax+Bx^2+4B+Cx+D\]

myininaya (myininaya):

\[x^3=Ax^3+Bx^2+x(4A+C)+(4B+D)\]

myininaya (myininaya):

\[=>A=1, B=0, 4A+C=0, 4B+D=0\]

myininaya (myininaya):

=>C=-4, D=0

myininaya (myininaya):

\[\int\limits_{}^{}(\frac{1x}{x^2+4}+\frac{-4x}{(x^2+4)^2}) dx\]

myininaya (myininaya):

this should be easy now for you just use a substitution like u=x^2+4 =>du=2x dx

OpenStudy (anonymous):

what happens if you try \[u=x^2+4\] \[du=2xdx\] \[x^2=u-4\] ? this is an honest question because i am not sure

OpenStudy (anonymous):

yea i got it from here thank you

OpenStudy (anonymous):

dont think substitution would work cuz the top is X^3

OpenStudy (anonymous):

maybe not but let me try anyway. i know we have \[x^3dx\] but if i use \[\frac{1}{2}du=xdx\] all i need in the top is \[x^2=u-4\]

myininaya (myininaya):

are you asking is there a way to avoid writing it by partial fractions?

OpenStudy (anonymous):

yes i am just wondering (not being a critic)

myininaya (myininaya):

you could use a trig substitution

OpenStudy (anonymous):

what about \[\frac{1}{2}\int\frac{u-4}{u^2}du\] do we get the same answer?

myininaya (myininaya):

yep that looks like it works too gj

OpenStudy (anonymous):

\[\frac{1}{2}[\int\frac{1}{u}du-\int\frac{4}{u^2}du]\] \[\frac{1}{2}\ln(u)+\frac{2}{u}\]

OpenStudy (anonymous):

\[\frac{1}{2}\ln(x^2+4)+\frac{2}{x^2+4}\] is that the same?

myininaya (myininaya):

i bet so i see nothing wrong with your way

myininaya (myininaya):

+C

myininaya (myininaya):

not they are certainly the same

myininaya (myininaya):

now*

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

now if i don't get some actual work done i am going to be in trouble

myininaya (myininaya):

yep i need to read a paper

OpenStudy (amistre64):

\[\int \frac{x^3}{(x^2+4)^2}dx\] \[\int x^2\frac{x}{(x^2+4)^2}dx\] \[\frac{1}{2}\int x^2\frac{2x}{(x^2+4)^2}dx\] \[\frac{1}{2}(-\frac{x^2}{x^2+4}-\int \frac{2x}{x^2+4}dx)\] \[-\frac{x^2}{2(x^2+4)}-\frac{1}{2}\int \frac{2x}{x^2+4}dx\] \[-\frac{x^2}{2(x^2+4)}-\frac{1}{2}ln|x^2+4|+C\] whered i mess up ?

OpenStudy (amistre64):

i dropped a negative ... i see that

OpenStudy (anonymous):

more than one way to skin a cat